Session: Sep 1997 (session 3)
Frequency: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Cambridge (Alastair Gunn at Fri Sep 12 10:10:58 METDST 1997) Failure
This observation was performed as an operational test as no Tsys logging information is available at present over the Cambridge link. Schedule was stopped at 17:08 UT to check antenna control problem.

Onsala (John Conway at Sat Sep 13 16:33:43 METDST 1997) Success
No problems

Effelsberg (dgraham at Mon Sep 15 10:51:23 METDST 1997) Failure
3C274 unsuitable, 20Jy over 150", but our beam is 43". The results are useless.

Onsala (John Conway at Tue Sep 16 01:03:27 METDST 1997) Minor failures
I disagree with some of Daves figures, Barrs et al and published VLA maps at 5 and 22GHz show the lobe structure contained within 40 arcsecond OK for everyone except Bonn. The larger halo contributes less tha 1 percent at greater 10GHz. I had understood from the scheduler that Bonn would not be taking part

Metsahovi (Kirsi Karlamaa \& Kaj Wiik at Tue Sep 16 13:21:39 METDST 1997) Success
Started late (1457) because of tape unloading and procedure problems. First about 50% cloud cover, some seemed to be quite wet. After about 1530 clear sky in the direction of the source.

Noto (C.Trigilio at Mon Sep 22 16:30:34 METDST 1997) Success
The observations was OK. Check of gain curve was performed on DR21 (19 Jy) before the start of the session. Values of the actual caltemps already given.

Medicina (orfei at Sat Sep 27 10:20:33 METDST 1997) Success
Sorry for the delay of the report. The observation was done with local procedure using onoff program. No known problem encountered but the data are not yet processed.